# On the Effectiveness of Large Language Models in Domain-Specific Code Generation

### v2

@@ Abstract.
conference: Conference on Software Engineering; Dec 2023;
@@ 1. Introduction
Large language models (LLMs) trained with publicly available code, such as ChatGPT, have demonstrated remarkable proficiency in various coding tasks, including code generation (Chen et al., 2021; Nijkamp et al., 2022), software testing (Siddiq et al., 2023), and program repair (Cao et al., 2023).
Despite their remarkable performance, current LLMs for code generation heavily rely on enormous training data to acquire a broad spectrum of open-domain knowledge. Notably, the evaluation of present LLMs typically revolves around open-domain benchmarks like HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), which primarily consist of programming contests (e.g., sorting, dynamic programming), and while they showcase the capabilities of LLMs in certain aspects, they do not fully represent the intricacies and challenges associated with real-world code generation scenarios.
@@
We first investigate the abilities of several LLMs (ChatGPT (AI, 2023), CodeGen (Nijkamp et al., 2022), and PolyCoder (Xu et al., 2022)) in specific domains by comparing their performance on general-purpose code corpora and domain-specific code corpora. Our analysis reveals that although LLMs have made remarkable advancements in generating code for open-domain applications, their performance degrades sharply when applied to specific domains. For ChatGPT, the CodeBLEU score drops by 51.48% on average. We particularly notice that this is often caused by a lack of domain knowledge, particularly the misuse of third-party libraries.
@@
We then investigate how to effectively prompt LLMs using domain knowledge. We designed several basic knowledge-based prompts and use them to elicit ChatGPT, the most popular LLM for code generation. We experimented with different combinations of the basic prompts to study how different types of prompts influence the performance of domain-specific code generation. We observe result improvement consistently on all library-specific datasets when using knowledge-enhanced prompts such as API sequences and docstrings compared to using the plain function signature prompt.
@@
Based on these findings, we further investigate how to efficiently integrate domain knowledge into the code generation process of LLMs. We experiment with three strategies, including 1) inquiring an external knowledge GPT (denoted as kg-GPT), 2) Chain-of-though prompting (CoT-PT), and 3) Chain-of-thought fine-tuning (CoT-FT). In the kg-GPT strategy, we train an additional GPT model which predicts the API call sequence based on the function signature and prompts LLM to complete the function body by utilizing the API prediction as a prompt. In the chain-of-thought strategies, we conceptualize a domain function as a series of sequential states, where each state corresponds to a sub-activity related to an API call to a third-party library. We simulate a chain-of-thought process during code generation: at each step, the model generates a ‘knowledge state’ suggesting relevant APIs and a ‘task state’ describing the intended action to be performed. A knowledge state is predicted based on the history of preceding steps, acting as a guide for the code generation at the current step and augmenting domain-specific knowledge. The process continues until the entire programming task is completed. We implement two distinct variants of the chain-of-thought strategies, encompassing both the zero-shot and fine-tuning paradigms. Concerning zero-shot code generation (CoT-PT), we sequentially predict APIs using kg-GPT and leverage each API prediction to prompt subsequent states in the sequence. In the fine-tuning setting (CoT-FT), we enrich each of the primary training functions by injecting pre-processed knowledge states. These enriched functions are then used to fine-tune LLMs in an end-to-end manner.
@@
We apply these strategies on PolyCoder-2.7B and StarCoder, under both zero-shot and fine-tuning paradigms. Our experimental results show that all strategies lead to improvement in the effectiveness of domain-specific code generation under certain settings. The results also show that there is still ample room for further improvement, based on which we suggest possible future works.
@@
  • An investigation of the effectiveness of various prompts for LLMs to generate domain-specific code, including a comparison of three strategies to integrate domain knowledge in the code generation process.
@@ 2.1. Code Generation
Code generation aims to automatically synthesize programs conditioned on programming language context and/or natural language descriptions (Chen et al., 2021; Xu et al., 2022; Black et al., 2022; Yusuf et al., 2022a), therefore increasing productivity for developers. Recently, LLMs for code generation have gained significant popularity due to their impressive performance. The state-of-art models are based on GPT, due to their auto-regressive nature, which is suitable for generative tasks where the model predicts the next tokens given the previous context.
@@ 2.2. Prompting Large Language Models
The use of prompting techniques has also been extended to code generation (Wang et al., 2022; Liu et al., 2023a; Choi and Lee, 2023; Cheng et al., 2023). Some researchers optimize the prompting patterns. Liu et al. (Liu et al., 2023a) and White et al. (White et al., 2023) investigated various prompt templates for ChatGPT. Some works enhance prompting by incorporating contextual information. For example, Liu et al. (Liu et al., 2023a) added to the prompt optimization template extra context such as Java Class information. Li et al. (Li et al., 2023b) introduced test cases and API information to retrieval-based prompts.
@@ 2.3. Code Generation with Domain Libraries
Recently, code generation utilizing domain libraries has attracted research attention and there have been numerous works focused on incorporating knowledge into code generation models. For example, Zan et al. (Zan et al., 2022a) propose a retrieval-then-coding framework where an API retriever first identifies useful APIs and an API Coder generates code using these APIs. Shrivastava et al. (Shrivastava et al., 2023) introduce domain-specific knowledge in the prompt design process. They present the Repo-Level Prompt Generator framework, which generates example-specific prompts for LLMs. This framework incorporates context from entire repositories. Liu et al. (Liu et al., 2023b) introduce CodeGen4Libs, a technique specifically designed for library-oriented code generation. CodeGen4Libs first generates import statements and then concrete code. Zan et al. (Zan et al., 2022b) leverage the fact that library-oriented code snippets often share similar code sketches. They propose a strategy involving a sketcher and a generator trained on unlabelled data. The sketcher generates library-oriented code sketches, then the generator predicts code snippets based on them.
@@ 3.2. Evaluated Language Models
ChatGPT (AI, 2023) has the ability to solve coding problems in the dialogue format, supported by a family of backbone models. We chose the most widely-used version gpt-3.5-turbo to experiment with. The model has around 154B parameters. Since it is not open-sourced, we connect to the model through the official API provided by OpenAI [^2] ChatGPT: https://platform.openai.com/docs/api-reference once at a time, then retrieve the code segment from its response text.
CodeGen (Nijkamp et al., 2022) is a family of open-source models for program synthesis. We chose the version CodeGen-Multi 2B, which is pre-trained on a multi-lingual dataset BIGQUERY (Google, 2021) (C, C++, Go, Java, JavaScript, and Python) and has a size of 2B. We use the released pre-trained checkpoints [^3] CodeGen: https://huggingface.co/Salesforce/codegen-2B-multi to conduct generation on each prompt.
PolyCoder (Xu et al., 2022): PolyCoder is a well-established open-source model based on GPT-2 and trained on GitHub code, which demonstrates competitive performance to GPT-3 counterparts. We chose the 2.7B version to experiment with. We use the released pre-trained checkpoint [^4] PolyCoder: https://huggingface.co/NinedayWang/PolyCoder-2.7B to conduct generation on each prompt.
@@ 3.3. Evaluation Metrics
We did not adopt the widely used pass@k (the passing rate of generated code on test cases) (Chen et al., 2021) for evaluation. The rationale behind this choice is rooted in the distinctive nature of the domain-specific functions we investigate. Unlike programming contests in existing benchmarks (Chen et al., 2021; Nijkamp et al., 2022), domain-specific functions tend to be notably intricate, often interrelated with multiple functions, designed for specific platforms, or reliant on a lot of third-party libraries. It is too rigorous to have all of them executable, even with powerful LLMs. For example, our preliminary experiments indicate that ChatGPT achieves a 0% success rate out of 200 samples when attempting to pass these test cases. Furthermore, testing them necessitates the labor-intensive task of manually crafting hundreds of test cases, thoroughly constructing multiple functions, and configuring reliant libraries and platforms, which is an impractical endeavor. Given these inherent complexities and limitations, conducting evaluations based on real executions is infeasible.
Therefore, in our study, we have chosen to utilize other popular metrics such as BLEU and CodeBLEU. BLEU (Papineni et al., 2002) is a widely used measurement for sequence similarity by calculating n-gram matches. CodeBLEU (Ren et al., 2020) is a metric based on BLEU, adjusted for code evaluation, which introduces syntactic and semantic similarities into comparison. We compute both metrics using the scripts provided by the CodeXGLUE paper [^5] https://github.com/microsoft/CodeXGLUE/tree/main/Code-Code/code-to-code-trans/evaluator/CodeBLEU.
@@
Dataset ChatGPT-3.5-154B CodeGen-2B-multi PolyCoder-2.7B
@@
 HumanEval 39.00 30.05 22.50 16.82 22.22 13.81
 CodeSearchNet (Golang) 11.16 25.29 14.61 19.68 15.97 19.56
@@
 Gin 5.81 17.73 7.73 16.09 8.13 17.05
 Prometheus 1.75 12.17 1.80 13.16 1.77 12.11
 gRPC-go 2.53 12.15 35.43 43.62 55.36 57.52
 Unreal Engine 5.22 10.57 5.82 11.30 0.94 9.87
 cocos2d-x 3.71 12.92 16.55 28.35 13.83 23.83
 bgfx 0.86 8.09 0.81 7.41 0.76 7.16
@@ 3.4.1. Methodology
In this RQ, we investigate the capabilities of LLMs in specific domains. We focus on the code completion scenario, wherein an LLM is presented with a function signature as a prompt, and its objective is to effectively generate the appropriate completion for the corresponding function body.
We sampled 500 functions from each library-specific dataset we collected (Section 3.1) and asked the LLM to complete the function body. We compare the model-generated code against the ground-truth one in the original function and compute the evaluation metrics.
@@ 3.4.2. Quantitative Results
We have observed that libraries that are not popular (such as Prometheus) tend to experience a more significant performance decline compared to others. This is likely due to the fact that there is less open-source code that utilizes the domain-specific library available for LLM training. On the contrary, gRPC-go and cocos2d-x exhibit exceptionally good performance in code-oriented LLMs. Upon manual examination, we observe monotonous code patterns in functions that utilize these two libraries, such as gRPC server handler functions and cocos2d-x interface functions from C++ to Lua. Such repetitive knowledge is more easily learned by the models.
@@ 3.5.2. Results
Comparing the results of basic prompts, API sequence prompts lift the result significantly, while library import prompts and docstring prompts show relatively smaller increases. The API name sequences are closer related to the code content and have a more direct effect on code generation. In contrast, the library import prompts do carry concise and useful information, but in meager quantity. The API docstrings contain implicit knowledge in natural language, different from the target programming language, which could be indirect.
Comparing all knowledge-enhanced prompts, using API sequence and docstring as prompt reaches the highest performance gain, suggesting that these two types of knowledge do have a complementary effect when combined together.
@@ 4.2.1. Programming as a chain-of-thought process
Like reasoning, programming is also a task that requires careful logical thinking (Yusuf et al., 2022b). When faced with a programming task to implement a complex requirement, one often thinks in a chain-of-thought manner (Cheng et al., 2023). This is natural since a complete code segment (e.g., a function) can usually be decomposed into multiple sub-activities (Sun et al., 2019), each containing one or a few statements. We refer to a sub-activity as a step in the coding process. At each step, one decides what action should be taken based on former contexts and how it should be implemented. For domain-specific code, this decomposition of activities is much related to API calls of third-party libraries (Nguyen et al., 2021, 2020). An API acts like an intermediary, providing developers with functionality implemented by a third-party library. One statement calling an API can be regarded as a separable sub-activity.
@@
Figure 3 shows an example of a function generation process in a chain-of-thought manner. We take a function signature as input, where the function name HandleGetRequest indicates that the requirement is to handle GET requests, and the parameter c *gin.context points to the Gin library. Based on the input, the model generates code step by step. At step 0, it first predicts the API state gin.Context.Query and the task state get the request parameter. The states suggest that the following code should perform the task of retrieving the request parameter using the API gin.Context.Query. Then conditioned on the code history (which is the function signature for step 0) and the current state, the model generates code for step 0, param := c.Query("param"). Similarly, the model continues to recursively predict code with the states at each step, until the end of the function is reached. If no API is required, we mark it as an empty state and simply skip the API prediction for this step.
@@ 4.2.2. Overall Design
Figure 2 (b) illustrates the design of this strategy. It consists of two modules, one for prompt generation and the other for code generation. The prompt generator is realized with an external knowledge enquiring model kg-GPT as shown in Figure 2 (a), it predicts the next possible API based on the API state history or input function signature. For each step, the prompt-generator predicts an API state based on the state history, the generated API state is combined with the code history and used as a prompt for the LLM, which then predicts code for the current step.

### v3

@@ Abstract.
conference: Conference on Software Engineering; May 2024;
@@ 1. Introduction
Large language models (LLMs), such as ChatGPT, have demonstrated remarkable proficiency in various coding tasks (Nascimento et al., 2023; Hansson and Ellréus, 2023), including code generation (Chen et al., 2021; Nijkamp et al., 2022), software testing (Siddiq et al., 2023), program repair (Cao et al., 2023), and code refinement (Guo et al., 2024; DePalma et al., 2024).
Despite their remarkable performance, current LLMs for code generation heavily rely on enormous training data to acquire a broad spectrum of open-domain knowledge. Notably, the evaluation of present LLMs typically revolves around open-domain benchmarks like HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), which primarily consist of programming contests (e.g., sorting, dynamic programming), and while they showcase the capabilities of LLMs in certain aspects, they do not fully represent the intricacies and challenges associated with real-world code generation scenarios (Liu et al., 2024; Jin et al., 2024).
@@
We first investigate the abilities of several LLMs (ChatGPT (AI, 2023), PolyCoder (Xu et al., 2022), and CodeLlama (Roziere et al., 2023)) in specific domains by comparing their performance on general-purpose code corpora and domain-specific code corpora. Our analysis reveals that although LLMs have made remarkable advancements in generating code for open-domain applications, their performance degrades sharply when applied to specific domains. For ChatGPT, the CodeBLEU score drops by 51.48% on average. We particularly notice that this is often caused by a lack of domain knowledge, particularly the misuse of third-party libraries.
@@
We then investigate how to effectively prompt LLMs using domain knowledge. We design several basic knowledge-based prompts and use them to elicit ChatGPT, the most popular LLM for code generation. We experimented with different combinations of the basic prompts to study how different types of prompts influence the performance of domain-specific code generation. We observe result improvement consistently on all library-specific datasets when using knowledge-enhanced prompts such as API sequences and docstrings compared to using the plain function signature prompt.
@@
Based on these findings, we propose a new method called DomCoder for domain-specific code generation. DomCoder integrates domain knowledge into the code generation process of LLMs through three strategies, including 1) inquiring an external knowledge GPT (denoted as kg-GPT), 2) Chain-of-though prompting (CoT-PT), and 3) Chain-of-thought fine-tuning (CoT-FT). Kg-GPT trains an external API enquirer based on GPT. Given a function signature, it predicts the API call sequence and uses it to prompt LLM to complete the function body. In the chain-of-thought strategies, we conceptualize a domain function as a series of sequential states, where each state corresponds to a sub-activity related to an API call to a third-party library. We simulate a chain-of-thought process during code generation: at each step, the model generates a ‘knowledge state’ suggesting relevant APIs and a ‘task state’ describing the intended action to be performed. A knowledge state is predicted based on the history of preceding steps, acting as a guide for the code generation at the current step and augmenting domain-specific knowledge. The process continues until the entire programming task is completed. We implement two distinct variants of the chain-of-thought strategies, encompassing both the zero-shot and fine-tuning paradigms. Concerning zero-shot code generation (CoT-PT), we sequentially predict APIs using kg-GPT and leverage each API prediction to prompt subsequent states in the sequence. In the fine-tuning setting (CoT-FT), we enrich each of the primary training functions by injecting pre-processed knowledge states. These enriched functions are then used to fine-tune LLMs in an end-to-end fashion.
@@
We apply these strategies on the state-of-the-art code LLMs such as PolyCoder (Xu et al., 2022), StarCoder (Li et al., 2023a), and CodeLlama (Roziere et al., 2023) under both zero-shot and fine-tuning paradigms. Our experimental results show that all strategies of DomCoder improve the effectiveness of domain-specific code generation under certain settings.
@@
  • An investigation of the effectiveness of various prompts for LLMs to generate domain-specific code.
  • A new approach to integrate domain knowledge into LLMs for code generation.
@@ 2.1. Code Generation
Code generation aims to automatically synthesize programs conditioned on programming language context and/or natural language descriptions (Chen et al., 2021; Xu et al., 2022; Black et al., 2022), therefore increasing productivity for developers. Recently, LLMs for code generation have gained significant popularity due to their impressive performance. The state-of-art models are based on GPT, due to their auto-regressive nature, which is suitable for generative tasks where the model predicts the next tokens given the previous context.
@@ 2.2. Prompting Large Language Models
The use of prompting techniques has also been extended to code generation (Wang et al., 2022; Liu et al., 2023a; Choi and Lee, 2023; Cheng et al., 2023). Some researchers optimize the prompting patterns. Liu et al. (Liu et al., 2023a) and White et al. (White et al., 2023) investigated various prompt templates for ChatGPT. Some works enhance prompting by incorporating contextual information. For example, Liu et al. (Liu et al., 2023a) added to the prompt optimization template extra context such as Java Class information. Li et al. (Li et al., 2023c) introduced test cases and API information to retrieval-based prompts.
@@ 2.3. Code Generation with Domain Libraries
Recently, code generation utilizing domain libraries has attracted research attention and there have been numerous works focused on incorporating knowledge into code generation models (Zan et al., 2022a; Liu et al., 2023b). For example, Zan et al. (Zan et al., 2022a) propose a retrieval-then-coding framework where an API retriever first identifies useful APIs and an API Coder generates code using these APIs. Shrivastava et al. (Shrivastava et al., 2023) introduce domain-specific knowledge in the prompt design process. They present the Repo-Level Prompt Generator framework, which generates example-specific prompts for LLMs. This framework incorporates context from entire repositories. Liu et al. (Liu et al., 2023b) introduce CodeGen4Libs, a technique specifically designed for library-oriented code generation. CodeGen4Libs first generates import statements and then concrete code. Zan et al. (Zan et al., 2022b) leverage the fact that library-oriented code snippets often share similar code sketches. They propose a strategy involving a sketcher and a generator trained on unlabelled data. The sketcher generates library-oriented code sketches, then the generator predicts code snippets based on them.
@@ 3.2. Evaluated Language Models
ChatGPT (AI, 2023) is a popular LLM that performs generation tasks in the dialogue format. The model also has demonstrated superb ability in coding. ChatGPT involves a family of backbone models. We chose the most widely-used version gpt-3.5-turbo to experiment with. The model has around 154B parameters. Since it is not open-sourced, we connect to the model through the official API provided by OpenAI [^2] ChatGPT: https://platform.openai.com/docs/api-reference once at a time, then retrieve the code segment from its response text.
PolyCoder (Xu et al., 2022): PolyCoder is a well-established open-source model based on GPT-2 and trained on GitHub code. The model demonstrates competitive performance to GPT-3 counterparts. We use the released pre-trained checkpoint PolyCoder-7B [^3] PolyCoder: https://huggingface.co/NinedayWang/PolyCoder-2.7B.
CodeLlama (Roziere et al., 2023) is the state-of-the-art language model for code generation. CodeLlama is built upon Llama 2, by further training on code-specific datasets. The model involves four sizes with 7B, 13B, 34B, and 70B parameters respectively. We use the 7B model, which can be served on a single GPU while showing efficient and accurate performance on code completion.
@@ 3.3. Evaluation Metrics
We did not adopt the widely used pass@k (the passing rate of generated code on test cases) (Chen et al., 2021) for evaluation. The rationale behind this choice is rooted in the distinctive nature of the domain-specific functions we investigate. Unlike programming contests in existing benchmarks (Chen et al., 2021; Nijkamp et al., 2022), domain-specific functions tend to be notably intricate, often interrelated with multiple functions, designed for specific platforms, or reliant on a lot of third-party libraries. It is too rigorous to have all of them executable, even with powerful LLMs. For example, our preliminary experiments indicate that even ChatGPT achieves a 0% pass rate out of 200 samples when attempting to pass their test cases. Furthermore, testing them necessitates the labor-intensive task of manually crafting hundreds of test cases, thoroughly constructing multiple functions, and configuring reliant libraries and platforms, which is an impractical endeavor. Given these inherent complexities and limitations, conducting evaluations based on real executions is infeasible.
Therefore, in our study, we have chosen to utilize other popular metrics such as BLEU and CodeBLEU, which have also been employed by related works on code generation (Li et al., 2023b; Liu et al., 2023b). BLEU (Papineni et al., 2002) measures sequence similarity by calculating their n-gram matches. CodeBLEU (Ren et al., 2020) is a metric based on BLEU, adjusted for code evaluation, which introduces syntactic and semantic similarities into comparison. We compute both metrics using the scripts provided by the CodeXGLUE paper [^4] https://github.com/microsoft/CodeXGLUE/tree/main/Code-Code/code-to-code-trans/evaluator/CodeBLEU.
@@
Dataset ChatGPT-3.5-154B CodeLlama-7B PolyCoder-2.7B
@@
HumanEval 39.00 30.05 18.13 17.99 22.22 13.81
CodeSearchNet (Golang) 11.16 25.29 16.14 19.63 15.97 19.56
@@
Gin 5.81 17.73 8.47 18.07 8.13 17.05
Prometheus 1.75 12.17 1.79 13.13 1.77 12.11
gRPC-go 2.53 12.15 57.61 60.39 55.36 57.52
Unreal Engine 5.22 10.57 0.73 10.21 0.94 9.87
cocos2d-x 3.71 12.92 16.76 25.87 13.83 23.83
bgfx 0.86 8.09 0.55 7.72 0.76 7.16
@@ 3.4.1. Methodology
In this RQ, we investigate the capabilities of LLMs in specific domains. We focus on the code completion scenario, wherein an LLM is presented with a function signature as a prompt, and its objective is to complete the function body. Previous work has indicated that a good function signature already provides enough information to illustrate its functionality (Ding et al., [n. d.]). Another reason for not including NL requirements in the prompt is that the domain-specific code we collected contains only a small portion ($<$5%) of code that contains NL requirements.
We sampled 500 functions from each library-specific dataset we collected (Section 3.1) for the completion. We compare the model-generated code against the ground-truth one in the original function and compute the evaluation metrics.
@@ 3.4.2. Quantitative Results
We have noted that non-popular libraries (such as Prometheus) tend to experience a more significant performance decline compared to others. This is likely due to the fact that there is less open-source code that utilizes the domain-specific library available for LLM training. On the contrary, gRPC-go and cocos2d-x exhibit exceptionally high performance in code-oriented LLMs. Upon manual examination, we observe monotonous code patterns in functions that utilize these two libraries, such as gRPC server handler functions and cocos2d-x interface functions from C++ to Lua. Such repetitive knowledge is more easily learned by the models.
@@ 3.5.2. Results
Comparing the results of basic prompts, API sequence demonstrates a significant improvement, whereas library import and docstring exhibit relatively modest enhancements. The API name sequences are closer related to the code content and have a more direct effect on code generation. In contrast, the library import prompts do carry concise and useful information, but in meager quantity. The API docstrings contain implicit knowledge in natural language, different from the target programming language, which could be indirect. Furthermore, we observed that the presence of docstrings alongside APIs does not consistently enhance performance when compared to using APIs alone, despite occasional cases where such combined prompts achieve the highest scores. This discrepancy may be attributed to the long text of the docstrings, which sometimes hinders the comprehension of APIs when incorporated into the prompt.
@@ 4.2.1. Programming as a Chain-of-thought Process
Like reasoning, programming is also a task that requires careful logical thinking (Yusuf et al., 2022). When faced with a programming task to implement a complex requirement, one often thinks in a chain-of-thought manner (Cheng et al., 2023). This is natural since a complete code segment (e.g., a function) can usually be decomposed into multiple sub-activities (Sun et al., 2019), each containing one or a few statements. We refer to a sub-activity as a step in the coding process. At each step, one decides what action should be taken based on former contexts and how it should be implemented. For domain-specific code, this decomposition of activities is much related to API calls of third-party libraries (Nguyen et al., 2021, 2020). An API acts like an intermediary, providing developers with functionality implemented by a third-party library. One statement calling an API can be regarded as a separable sub-activity.
@@
Figure 3 shows an example of a function generation process in a chain-of-thought manner. We take a function signature as input, where the function name HandleGetRequest indicates that the requirement is to handle GET requests, and the parameter c *gin.context points to the Gin library. Based on the input, the model generates code step by step. At step 0, it first predicts the API state gin.Context.Query and the task state get the request parameter. The states suggest that the following code should perform the task of retrieving the request parameter using the API gin.Context.Query. Then conditioned on the code history (which is the function signature for step 0) and the current state, the model generates code for step 0, param := c.Query(‘‘param"). Similarly, the model continues to recursively predict code with the states at each step, until the end of the function is reached. If no API is required, we mark it as an empty state and simply skip the API prediction for this step.
@@ 4.2.2. Overall Design
Figure 2 (b) illustrates the design of this strategy. It consists of two modules, one for prompt generation and the other for code generation. The prompt generator is realized with an external knowledge enquiring model kg-GPT as shown in Figure 2 (a), it predicts the next possible API based on the API state history or input function signature. For each step, the prompt-generator predicts an API state based on the state history. The corresponding API docstring is appended as the task state. The generated API and task states are combined with the code history and used as a prompt for the LLM, which then predicts code for the current step.